OpenStudy (didee):

Express sin 4 "theta" in terms of powers of sin "theta" and cos "theta". Please explain in detail, thanks

OpenStudy (callisto):

Is the question\[sin4\theta\]

OpenStudy (anonymous):

Use the formulae \[ \sin ( 2 A) = 2 \sin(A) \cos(A)\\ \cos(2 A)= \cos^2(A) - \sin^2(A) \] to get what you want.

OpenStudy (anonymous):

Your final answer should look like \[ \sin (4 \theta )=4 \sin (\theta ) \cos ^3(\theta )-4 \sin ^3(\theta ) \cos (\theta ) \] but you should do all the steps from the two formulae I wrote above.

OpenStudy (didee):

Hi, sorry for my late reply and thanks for your feedback. I'm clueless when it comes to this so can you give a bit more clarity?